Kevin used his credit card to pay $2544 for a holiday. The interest rate is 18.75 compounded daily. Kevin plans too make a monthly payment of $200.

when will kevin have paid off the Balance in full?


1
Expert's answer
2021-03-31T13:44:16-0400

As Kevin plans to make a monthly payment, a month is the period for calculations.

Present value PV = $2544

Periodic payment PMT = $200

Rate of interest per period

"i = \frac{18.75}{12}\%=1.5625 \% = 0.015625

PV = PMT( \frac{1-(1+i)^{-n}}{i} ) \\

2544 = 200( \frac{1-(1+0.015625)^{-n}}{0.015625} ) \\

1-(1.015625)^{-n} = 0.19875 \\

(1.015625)^{-n} = 1-0.19875 \\

(1.015625)^{-n} = 0.80125"

Taking logarithm of both sides, we get

"-nlog1.015625=log0.80125 \\

n = -\frac{log0.80125}{log1.015625} \\

n = 14.29"

Answer: 14 months and 9 days.